a cylinder has a radius of 15cm and a surface area of 1,380 cm2 . what is the height
Answer by lwsshak3(11628)

a cylinder has a radius of 15cm and a surface area of 1,380 cm2 . what is the height
***
let h=height of cylinder
h*circumference=area
h*2π*15=1380
h=1380/30π=46/π=14.64
height of cylinder=14.64 cm